# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v2

@@ 1 Introduction
In recent developments, graph neural network (GNN) techniques have been employed to analyse brain networks characterised by pairwise interactions. The utilisation of these methodologies enables the extraction of potential topological features, which are instrumental in the diagnosis of brain disorders (Luo et al., 2024). A number of GNN-based models have been proposed for brain networks, including GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024), and ASD-HNet (Luo et al., 2025). These models have demonstrated the potential to achieve favourable diagnostic performance. It is widely accepted that GNN-based models utilise a message passing mechanism, whereby the embedding of a brain ROI is updated by aggregating information from its neighbouring ROIs. This process facilitates the learning of a discriminative graph-level representation of the brain connectivity network. However, these methods are limited by the underlying assumption that interactions between nodes are strictly two-way. Recent studies have indicated a growing body of evidence that suggests the complexity of brain-region interactions extends far beyond the established pairwise connections. There is now a growing consensus that widespread co-fluctuations occur in groups of nodes that evolve over time (Battiston et al., 2020; Chelaru et al., 2021). Consequently, the limited expressiveness of traditional GNNs prevents them from capturing higher-order interactions in brain networks, and the topological features they extract are typically regarded as lower-order.
@@
A number of studies have concentrated on the representation of higher-order interactions (HOIs) in brain networks using more sophisticated models, with a particular focus on hypergraphs. This approach has been shown to enhance diagnostic performance for brain diseases, as evidenced by citations in the literature (Jie et al., 2016; Xiao et al., 2019; Wang et al., 2022; Hao et al., 2023; Wang et al., 2024a). Despite the encouraging outcomes demonstrated by HGNN-based models in characterising HOIs through hyperedges generated via k-nearest neighbour or k-hop neighbourhoods, this node-centric construction scheme appears to be incompatible with the concept of group dependence. This is due to its inability to capture authentic simultaneous interactions among groups of ROIs. Furthermore, it has been demonstrated that higher-order features extracted by HGNN-based models can obscure their relationship to specific topological or neurobiological phenomena (Kim et al., 2024; Su et al., 2024).
@@
In order to extract higher-order topological features that are more readily interpretable for the purpose of brain-disorder diagnosis, recent studies have adopted persistent homology (PH), a topological approach capable of reconstructing HOI structures and delivering state-of-the-art performance in characterising brain topological profiles (Hyekyoung et al., 2011; Sizemore et al., 2018). However, the majority of investigations into higher-order structures in brain networks have focused on 0-dimensional (connected components) and 1-dimensional (cycles) topological profiles formed by nodes and edges. These dimensions do not link HOIs in the brain to higher-order organisations that would provide a truly higher-dimensional perspective (Talesh Jafadideh and Mohammadzadeh Asl, 2022; Bian et al., 2024; Bhattacharya et al., 2025). In order to address the aforementioned limitation, researchers have recently employed synchronisation phenomena to construct more accurate simplicial complexes for modelling HOIs. Furthermore, they have utilised PH to capture 1-dimensional cycles that reflect triplet interactions (Santoro et al., 2024). Comprehensive analyses demonstrate that methods incorporating inferred higher-order interactions among three ROIs outperform traditional pairwise approaches, thereby offering new insights into the higher-order organisation of fMRI time-series data. Nonetheless, the extant evidence indicates that triplet interactions can be decomposed into linear combinations of pairwise interactions, provided that said interactions are linearly decomposable. This finding suggests that some triplets may not in fact represent genuine higher-order phenomena (Delabays et al., 2025; Neuhäuser et al., 2020). Consequently, consideration of higher-order interactions and organisational structures may offer a more profound comprehension of brain function. Furthermore, extant measures of instantaneous co-fluctuation are contingent on extended Pearson correlation, a method that is deficient in temporal resolution when it comes to detecting spatiotemporal interactions among groups of regions (Shine et al., 2015). Finally, the role of signed HOIs, which have the potential to provide valuable diagnostic information in cases of neurological conditions, has been largely overlooked in these studies.
@@
In order to address the methodological challenges identified, a novel computational framework has been proposed. This framework, designated HOI-Brain (Higher-Order Interactions in Brain), facilitates a comprehensive analysis of HOIs in fMRI data 1. Firstly, a novel metric – Multiplication of Temporal Derivatives (MTD) – is introduced with the aim of quantifying dynamic functional co-fluctuations among groups of ROIs. MTD performs element-wise product calculations on the temporal derivatives of blood-oxygen-level-dependent (BOLD) signals, thereby yielding instantaneous co-fluctuation magnitudes for k-node interactions. This approach has been demonstrated to offer a more reliable means of identifying genuine higher-order neural interactions in comparison to Pearson’s correlation. Secondly, adopting a spatiotemporal analytic perspective, we propose a model for brain networks as weighted simplicial complexes. These complexes link instantaneous co-fluctuation of k-node interactions in the brain to the weight of the k-simplex. This facilitates the accommodation of more information about interactions among multiple regions of interest (ROIs) in comparison to pairwise networks. The following step involves the extraction of both positive and negative synergistic interactions from the weighted simplicial complexes in order to construct two signed weighted simplicial complexes, respectively. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer detailed information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. The present study explores the influence of signed higher-order interactions (HOIs) using the self-attention mechanism. The extraction of four types of higher-order topological features is achieved through the implementation of two distinct filtration processes, which are based on Persistent Homology theory. The first type of interaction signature is positive and negative quadruplet-level interaction signatures. These capture irreducible higher-order neural coordination patterns. The second type of interaction signature is positive and negative two-dimensional void descriptors. These characterise the intrinsic geometric organisation of neural activity from a higher-dimensional manifold perspective. Thirdly, given the significance of both lower-order interactions and higher-order interactions in brain network analysis tasks, a multi-channel brain network Transformer is developed to synergistically integrate lower-order edge features with the above four types of higher-order topological invariants. Specifically, higher-order information is injected into a multi-channel Transformer, and an orthonormal clustering readout operation is employed. This operation is based on self-supervised soft clustering and orthonormal projection, and it is used to learn distinguishable cluster-aware lower- and higher-order node embeddings. Subsequently, an attention mechanism is employed to facilitate the adaptive fusion of features encoded by disparate channels. In conclusion, the Multi-Layer Perceptron (MLP) is employed as the prediction head to categorise the various levels of topological features that have been extracted from the multi-channel brain network Transformer.
@@
    To the best of our knowledge, this is the first attempt to distinguish between positively and negatively synergistic higher-order interactions, which may offer valuable insights into the complex coordination and communication within the brain, thereby improving the effectiveness of brain disease diagnosis.
@@ 2.1 Graph models based on deep learning
A significant number of graph models are then applied to brain networks for the purpose of brain disorder diagnosis, with the objective being to understand and analyse the elements and interactions of neurological systems from a network perspective. Graph neural networks (GNNs) have been demonstrated to be particularly effective in modelling brain connectomes, due to their ability to effectively capture subtle, latent representations and nonlinear relationships in graph-structured data. A number of GNN-based models, specifically designed for brain networks, have been proposed, and these have achieved a promising diagnostic performance. For instance, GroupINN (Yan et al., 2019) incorporates the concept of node grouping into the neural network and designs a random-walk-based variant of graph convolutional layer. BrainGNN (Li et al., 2021) has developed a novel ROI-aware graph convolutional (Ra-GConv) layer that utilises the topological and functional information of fMRI. FBNETGEN (Kan et al., 2022a) employs a task-aware GNN-based framework for fMRI analysis via functional brain network generation, which generates the brain connectivity matrices and predicts clinical outcomes simultaneously from fMRI BOLD signal series. BPI-GNN (Zheng et al., 2024) utilises the prototype learning method to analyse functional magnetic resonance images (fMRI). ASD-HNNet (Luo et al., 2025) is a hybrid neural network model for the identification of autism spectrum disorder (ASD). This model extracts features from functional brain networks at three distinct levels: local ROI, community, and global representation. However, these methodologies are fundamentally predicated on a two-way network to model pairwise interactions and mine low-order topological features for the diagnosis of brain disorders, overlooking the influence of HOIs in the brain.
@@
Furthermore, with the advent of sophisticated large language models, Transformer-based methodologies are being increasingly incorporated into brain graph structures. The Graph Transformer (Ying et al., 2021), a sophisticated machine learning algorithm, can be applied to brain networks to learn the strength of the connections between regions of interest (ROIs) across individuals. Subsequently, the Brain Network Transformer (BNT) (Kan et al., 2022b) employs the distinctive characteristics inherent in brain network data to maximise the efficacy of Transformer-based models for brain network analysis, thereby circumventing the necessity for time-consuming computations of eigenvalues or eigenvectors. TSEN (Hu et al., 2023) pioneered the incorporation of snowball graph convolution as position embedding within the Transformer structure, a methodology that has been proven to be both straightforward and efficacious in the context of capturing local patterns of brain activity in a natural manner. Long-range Brain Transformer (Yu et al., 2024) injects the long-range embeddings into a Transformer framework, integrating both short-range and long-range dependencies between regions of interest (ROIs) using the self-attention mechanism. It is evident that these works disregard substantial higher-order topological information among ROIs.
@@
A number of studies have been conducted on the construction of HOIs in the brain with a view to enhancing the performance of brain disorder diagnosis based on hypergraphs. For instance, a novel framework is proposed to estimate the hyper-connectivity network of brain functions for the diagnosis of brain disease (Jie et al., 2016). MHL-Hypergraph (Xiao et al., 2019) proposes a multi-hypergraph learning-based method to compute a unified hypergraph similarity matrix from multi-paradigm fMRI data to represent an FCN for each subject. An evolving hypergraph convolutional network (Wang et al., 2022) for the dynamic hyper-brain network is proposed, which adds the attention mechanism to further enhance the ability of representation learning. In order to take the temporal characteristics of longitudinal data into consideration, a weighted hypergraph convolution network (WHGCN) has been designed to utilise the internal correlations among different time points and to leverage higher-order relationships between subjects for the purpose of AD detection (Hao et al., 2023). CcSi-MHAHGEL (Wang et al., 2024a) pioneered a novel hypergraph convolution network framework for extracting multiatlas-based FCN embeddings, with the objective of facilitating multisite ASD identification. In this framework, hyperedge-aware HGCN was developed to capture complex high-order information in brain networks. The majority of these studies employ k-nearest neighbour or k-hop neighbour constructions in order to generate hypergraphs. However, it has been demonstrated that these constructions are inconsistent with the definition of group dependence in brain regions. Moreover, higher-order features extracted using hypergraph neural networks (HGNNs) are frequently regarded as being less interpretable.
@@ 2.2 Persistent Homology on brain connectome
Persistent homology (PH) is a widely utilised, effective algebraic topological tool for the analysis of the brain connectome, enabling the capture of more interpretable, higher-order topological features. In lieu of endeavouring to ascertain a solitary optimum threshold, researchers (Hyekyoung et al., 2011) propose the examination of the topological changes in the brain network as the threshold is increased continuously, based on persistent homology. Researchers (Sizemore et al., 2018) also highlight the importance of cliques and cavities in the human connectome, and locate topological cavities of different dimensions, around which information may flow in either diverging or converging patterns. In order to address the challenging issue of comparing functional connectivity networks (FCNs) across different spatiotemporal resolutions, researchers (Cassidy et al., 2018) have developed a novel network comparison framework based on persistent homology. The purpose of this framework is to observe the change of 0-dimensional homology groups. Furthermore, researchers (Bian et al., 2024) propose an adversarially trained persistent homology-based graph convolutional network (ATPGCN) to capture disease-specific brain connectome patterns for brain disorder diagnosis. Researchers (Bhattacharya et al., 2025) utilise the persistent topological features of connected components $H_{0}$ and loops $H_{1}$, which are quantified by two methods: The utilisation of Vietoris-Rips filtration and graph filtration is a pivotal diagnostic tool in the identification of brain disorders. Nevertheless, the aforementioned studies have overlooked the significance of two-dimensional topological profiles (i.e., voids) and have failed to establish a correlation between HOIs in the brain and these higher-order organisations. Recently, researchers (Santoro et al., 2024) have adopted certain methodologies grounded in the synchronisation phenomenon to construct more accurate simplicial complexes for modeling HOIs and use PH to capture 1-dimensional cycles related to triplet interactions. However, this can easily generate a large number of false positive higher-order interactions and overlook higher-order quadruplet interactions with adequate temporal resolution. Besides, the influence of signed HOIs was overlooked in these studies, which could offer a broader diagnostic view of neurological conditions.

### v3

@@ 1 Introduction
In recent developments, graph neural network (GNN) techniques have been employed to analyse brain networks characterised by pairwise interactions. The utilisation of these methodologies enables the extraction of potential topological features, which are instrumental in the diagnosis of brain disorders (Luo et al., 2024). A number of GNN-based models have been proposed for brain networks, including GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024), and ASD-HNet (Luo et al., 2025). These models have demonstrated the potential to achieve favourable diagnostic performance. It is widely accepted that GNN-based models utilise a message passing mechanism, whereby the embedding of a brain ROI is updated by aggregating information from its neighbouring ROIs. This process facilitates the learning of a discriminative graph-level representation of the brain connectivity network. However, these methods are limited by the underlying assumption that interactions between nodes are strictly two-way. Recent studies have indicated a growing body of evidence that suggests the complexity of brain-region interactions extends far beyond the established pairwise connections. There is now a growing consensus that widespread co-fluctuations occur in groups of nodes that evolve over time (Battiston et al., 2020; Chelaru et al., 2021). Consequently, the limited expressiveness of traditional GNNs prevents them from capturing higher-order interactions (HOIs) in brain networks, and the topological features they extract are typically regarded as lower-order.
@@
A number of studies have concentrated on the representation of HOIs in brain networks using more sophisticated models, with a particular focus on hypergraphs. This approach has been shown to enhance diagnostic performance for brain diseases, as evidenced by citations in the literature (Jie et al., 2016; Xiao et al., 2019; Wang et al., 2022; Hao et al., 2023; Wang et al., 2024a). Despite the encouraging outcomes demonstrated by HGNN-based models in characterising HOIs through hyperedges generated via k-nearest neighbour or k-hop neighbourhoods, this node-centric construction scheme appears to be incompatible with the concept of group dependence. This is due to its inability to capture authentic simultaneous interactions among groups of ROIs. Furthermore, it has been demonstrated that higher-order features extracted by HGNN-based models can obscure their relationship to specific topological or neurobiological phenomena (Kim et al., 2024; Su et al., 2024).
@@
In order to extract higher-order topological features that are more readily interpretable for the purpose of brain-disorder diagnosis, recent studies have adopted persistent homology (PH), a topological approach capable of reconstructing HOI structures and delivering state-of-the-art performance in characterising brain topological profiles (Hyekyoung et al., 2011; Sizemore et al., 2018). However, the majority of investigations into higher-order structures in brain networks have focused on 0-dimensional (connected components) and 1-dimensional (cycles) topological profiles formed by nodes and edges. These dimensions do not link HOIs in the brain to higher-order organisations that would provide a truly higher-dimensional perspective (Talesh Jafadideh and Mohammadzadeh Asl, 2022; Bian et al., 2024; Bhattacharya et al., 2025). In order to address the aforementioned limitation, researchers have recently employed synchronisation phenomena to construct more accurate simplicial complexes for modelling HOIs. Furthermore, they have utilised PH to capture 1-dimensional cycles that reflect triplet interactions (Santoro et al., 2024). Comprehensive analyses demonstrate that methods incorporating inferred HOIss among three ROIs outperform traditional pairwise approaches, thereby offering new insights into the higher-order organisation of fMRI time-series data. Nonetheless, the extant evidence indicates that triplet interactions can be decomposed into linear combinations of pairwise interactions, provided that said interactions are linearly decomposable. This finding suggests that some triplets may not in fact represent genuine higher-order phenomena (Delabays et al., 2025; Neuhäuser et al., 2020). Consequently, consideration of HOIs and organisational structures may offer a more profound comprehension of brain function. Furthermore, extant measures of instantaneous co-fluctuation are contingent on extended Pearson correlation, a method that is deficient in temporal resolution when it comes to detecting spatiotemporal interactions among groups of regions (Shine et al., 2015). Finally, the role of signed HOIs, which have the potential to provide valuable diagnostic information in cases of neurological conditions, has been largely overlooked in these studies.
@@
In order to address the methodological challenges identified, a novel computational framework has been proposed. This framework, designated HOI-Brain (Higher-Order Interactions in Brain), facilitates a comprehensive analysis of HOIs in fMRI data 1. Firstly, a novel metric – Multiplication of Temporal Derivatives (MTD) – is introduced with the aim of quantifying dynamic functional co-fluctuations among groups of ROIs. MTD performs element-wise product calculations on the temporal derivatives of blood-oxygen-level-dependent (BOLD) signals, thereby yielding instantaneous co-fluctuation magnitudes for k-node interactions. This approach has been demonstrated to offer a more reliable means of identifying genuine higher-order neural interactions in comparison to Pearson’s correlation. Secondly, adopting a spatiotemporal analytic perspective, we propose a model for brain networks as weighted simplicial complexes. These complexes link instantaneous co-fluctuation of k-node interactions in the brain to the weight of the k-simplex. This facilitates the accommodation of more information about interactions among multiple ROIs in comparison to pairwise networks. The following step involves the extraction of both positive and negative synergistic interactions from the weighted simplicial complexes in order to construct two signed weighted simplicial complexes, respectively. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer detailed information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. The present study explores the influence of signed HOIs using the self-attention mechanism. The extraction of four types of higher-order topological features is achieved through the implementation of two distinct filtration processes, which are based on Persistent Homology theory. The first type of interaction signature is positive and negative quadruplet-level interaction signatures. These capture irreducible higher-order neural coordination patterns. The second type of interaction signature is positive and negative two-dimensional void descriptors. These characterise the intrinsic geometric organisation of neural activity from a higher-dimensional manifold perspective. Thirdly, given the significance of both lower-order interactions (LOIs) and HOIs in brain network analysis tasks, a multi-channel brain network Transformer is developed to synergistically integrate lower-order edge features with the above four types of higher-order topological invariants. Specifically, higher-order information is injected into a multi-channel Transformer, and an orthonormal clustering readout operation is employed. This operation is based on self-supervised soft clustering and orthonormal projection, and it is used to learn distinguishable cluster-aware lower- and higher-order node embeddings. Subsequently, an attention mechanism is employed to facilitate the adaptive fusion of features encoded by disparate channels. In conclusion, the Multi-Layer Perceptron (MLP) is employed as the prediction head to categorise the various levels of topological features that have been extracted from the multi-channel brain network Transformer.
@@
    To the best of our knowledge, this is the first attempt to distinguish between positively and negatively synergistic HOIs, which may offer valuable insights into the complex coordination and communication within the brain, thereby improving the effectiveness of brain disease diagnosis.
@@ 2.1 Graph models based on deep learning
A significant number of graph models are then applied to brain networks for the purpose of brain disorder diagnosis, with the objective being to understand and analyse the elements and interactions of neurological systems from a network perspective. Graph neural networks (GNNs) have been demonstrated to be particularly effective in modelling brain connectomes, due to their ability to effectively capture subtle, latent representations and nonlinear relationships in graph-structured data. A number of GNN-based models, specifically designed for brain networks, have been proposed, and these have achieved a promising diagnostic performance. For instance, GroupINN (Yan et al., 2019) incorporates the concept of node grouping into the neural network and designs a random-walk-based variant of graph convolutional layer. BrainGNN (Li et al., 2021) has developed a novel ROI-aware graph convolutional (Ra-GConv) layer that utilises the topological and functional information of fMRI. FBNETGEN (Kan et al., 2022a) employs a task-aware GNN-based framework for fMRI analysis via functional brain network generation, which generates the brain connectivity matrices and predicts clinical outcomes simultaneously from fMRI BOLD signal series. BPI-GNN (Zheng et al., 2024) utilises the prototype learning method to analyse fMRI. ASD-HNNet (Luo et al., 2025) is a hybrid neural network model for the identification of autism spectrum disorder (ASD). This model extracts features from functional brain networks at three distinct levels: local ROI, community, and global representation. However, these methodologies are fundamentally predicated on a two-way network to model pairwise interactions and mine low-order topological features for the diagnosis of brain disorders, overlooking the influence of HOIs in the brain.
@@
Furthermore, with the advent of sophisticated large language models, Transformer-based methodologies are being increasingly incorporated into brain graph structures. The Graph Transformer (Ying et al., 2021), a sophisticated machine learning algorithm, can be applied to brain networks to learn the strength of the connections between ROIs across individuals. Subsequently, the Brain Network Transformer (BNT) (Kan et al., 2022b) employs the distinctive characteristics inherent in brain network data to maximise the efficacy of Transformer-based models for brain network analysis, thereby circumventing the necessity for time-consuming computations of eigenvalues or eigenvectors. TSEN (Hu et al., 2023) pioneered the incorporation of snowball graph convolution as position embedding within the Transformer structure, a methodology that has been proven to be both straightforward and efficacious in the context of capturing local patterns of brain activity in a natural manner. Long-range Brain Transformer (Yu et al., 2024) injects the long-range embeddings into a Transformer framework, integrating both short-range and long-range dependencies between ROIs using the self-attention mechanism. It is evident that these works disregard substantial higher-order topological information among ROIs.
@@
A number of studies have been conducted on the construction of HOIs in the brain with a view to enhancing the performance of brain disorder diagnosis based on hypergraphs. For instance, a novel framework is proposed to estimate the hyper-connectivity network of brain functions for the diagnosis of brain disease (Jie et al., 2016). MHL-Hypergraph (Xiao et al., 2019) proposes a multi-hypergraph learning-based method to compute a unified hypergraph similarity matrix from multi-paradigm fMRI data to represent an FCN for each subject. An evolving hypergraph convolutional network (Wang et al., 2022) for the dynamic hyper-brain network is proposed, which adds the attention mechanism to further enhance the ability of representation learning. In order to take the temporal characteristics of longitudinal data into consideration, a weighted hypergraph convolution network (WHGCN) has been designed to utilise the internal correlations among different time points and to leverage higher-order relationships between subjects for the purpose of AD detection (Hao et al., 2023). CcSi-MHAHGEL (Wang et al., 2024a) pioneered a novel hypergraph convolution network framework for extracting multiatlas-based FCN embeddings, with the objective of facilitating multisite ASD identification. In this framework, hyperedge-aware HGCN was developed to capture complex higher-order information in brain networks. The majority of these studies employ k-nearest neighbour or k-hop neighbour constructions in order to generate hypergraphs. However, it has been demonstrated that these constructions are inconsistent with the definition of group dependence in brain regions. Moreover, higher-order features extracted using hypergraph neural networks (HGNNs) are frequently regarded as being less interpretable.
@@ 2.2 Persistent Homology on brain connectome
Persistent homology (PH) is a widely utilised, effective algebraic topological tool for the analysis of the brain connectome, enabling the capture of more interpretable, higher-order topological features. In lieu of endeavouring to ascertain a solitary optimum threshold, researchers (Hyekyoung et al., 2011) propose the examination of the topological changes in the brain network as the threshold is increased continuously, based on persistent homology. Researchers (Sizemore et al., 2018) also highlight the importance of cliques and cavities in the human connectome, and locate topological cavities of different dimensions, around which information may flow in either diverging or converging patterns. In order to address the challenging issue of comparing functional connectivity networks (FCNs) across different spatiotemporal resolutions, researchers (Cassidy et al., 2018) have developed a novel network comparison framework based on persistent homology. The purpose of this framework is to observe the change of 0-dimensional homology groups. Furthermore, researchers (Bian et al., 2024) propose an adversarially trained persistent homology-based graph convolutional network (ATPGCN) to capture disease-specific brain connectome patterns for brain disorder diagnosis. Researchers (Bhattacharya et al., 2025) utilise the persistent topological features of connected components $H_{0}$ and loops $H_{1}$ for the identification of brain disorders, which are quantified by two methods: Vietoris-Rips filtration and graph filtration. Nevertheless, the aforementioned studies have overlooked the significance of two-dimensional topological profiles (i.e., voids) and have failed to establish a correlation between HOIs in the brain and these higher-order organisations. Recently, researchers (Santoro et al., 2024) have adopted certain methodologies grounded in the synchronisation phenomenon to construct more accurate simplicial complexes for modeling HOIs and use PH to capture 1-dimensional cycles related to triplet interactions. However, this can easily generate a large number of false positive HOIs and overlook higher-order quadruplet interactions with adequate temporal resolution. Besides, the influence of signed HOIs was overlooked in these studies, which could offer a broader diagnostic view of neurological conditions.